### v1

# Isocurvature Induced Gravitational Waves at Pulsar Timing Arrays

Abstract
[…]
1 Introduction
[…]
Gravitational waves (GWs) serve as another powerful probe of early-Universe physics. Because they free-stream, GWs can access information from much earlier times than the CMB photons. This prospect drives the continued, worldwide development of GW observatories designed to measure a broad range of frequencies. In particular, Pulsar Timing Arrays (PTA) provide a unique probe of GWs in the nano-Hertz frequency range. Several experiments such as NANOGrav [NANOGrav:2023gor], EPTA/InPTA [EPTA:2023fyk], PPTA [Reardon:2023gzh], CPTA [Xu:2023wog] have recently observed a stochastic GW background and derived constraints for new physics. One possible source of GW that might contribute to PTA observations is called scalar induced GWs (SIGW) [^1] See Ref. [Domenech:SIGWRe] for a review. [Matarrese:1992rp, Matarrese:1993zf, Matarrese:1997ay, Ananda:2006af, Baumann:2007zm]: primordial scalar perturbations can source GWs at the second order in the perturbation theory. This framework has been applied to adiabatic perturbations and a certain type of isocurvature in cold dark matter (CDM) [Domenech:2021and, Domenech:2023jve, Chen:2024twp, Yuan:2024qfz, He:2024luf, Luo:2025lgr, Han:2025wlo, Ali:2025xvj, Zeng:2025tno, Domenech:2025ffb, Papanikolaou:2025ddc, Garcia:2025yit, Yuan:2025seu]. In this work, we present a more comprehensive study of isocurvature induced GWs including four types of isocurvature: cold dark matter density isocurvature (CDI), baryon density isocurvature (BDI), neutrino density isocurvature (NDI) and free-streaming dark radiation density isocurvature (DRDI). We emphasize that, for the frequency range relevant for PTA observations, the modes entered the horizon before SM neutrino decoupling (when the temperature of the Universe is around MeV), and thus neutrinos are tightly coupled to the SM bath. Therefore, the initial conditions for NDI are qualitatively different from the standard ones where neutrinos are treated as free-streaming radiation. In this work, we derive novel initial conditions for all isocurvature modes with coupled neutrinos, and calculate the evolution of density and metric perturbations numerically. We find that the GW spectrum from NDI with coupled neutrinos is the same as that of CDI or BDI up to an overall coefficient, which leads to an interesting ‘conversion’ of isocurvature between radiation and matter sectors. Furthermore, DRDI results are dramatically different from those in the previously studied CDI case because of the presence of anisotropic stress. The size of DRDI induced GWs is significantly enhanced for high $k$ modes compared to that of CDI because the ratio of dark radiation (DR) density to the total bath is not suppressed at early times unlike the matter density. Therefore, PTA opens a new window to test BSM physics and places novel constraints via isocurvature induced GWs.
To derive general constraints that can be applied to a broad class of models, we use two parameterizations of the primordial isocurvature power spectrum: a delta function and a broken power law [Buckley:2025zgh]. Using the general formula for the source term of induced GWs, we obtain the spectrum of GWs and apply the data from NANOGrav collaboration to place constraints on the isocurvature power spectrum. Our results set stringent limits for wavenumbers around $10^{6}\,\text{Mpc}^{-1}$, which are complementary to other cosmological constraints on large scales.
This paper is organized as follows. In section 2 we present new isocurvature initial conditions before neutrino decoupling. We show a detailed calculation of GWs induced by general isocurvature perturbations in section 3. We derive constraints on the isocurvature power spectrum from PTA in section 4. We conclude in section 5.
2 Isocurvature Initial Conditions Before Neutrino Decoupling
[…]
To study perturbations in cosmology [Kodama:1986fg, Kodama:1986ud, Ma:1995ey], we start with the perturbed FRW metric in the conformal Newtonian gauge,
$ds^{2}=a^{2}(\tau)\left[-(1+2\Phi)d\tau^{2}+(\delta_{ij}-2\Psi\delta_{ij}+h_{ij})dx^{i}dx^{j}\right]~,$ (2.1)
where $a$ is the scale factor and $\tau$ is the conformal time. $\Phi$ and $\Psi$ are two scalar metric perturbations. $h_{ij}$ denotes the tensor perturbation that satisfies the transverse and traceless conditions: $\partial^{i}h_{ij}=0,h^{i}_{i}=0$.
We consider the early Universe is dominated by radiation and matter, and thus the total background energy density $\bar{\rho}$ is given as
$\bar{\rho}$ $=$ $\bar{\rho}_{r}+\bar{\rho}_{m}.$ (2.2)
Here $\bar{\rho}_{r/m}$ is the background energy density of radiation and matter, which satisfies the following evolution equations:
$\bar{\rho}_{r}^{\prime}+4\mathcal{H}\bar{\rho}_{r}=0~~;~~\bar{\rho}_{m}^{\prime}+3\mathcal{H}\bar{\rho}_{m}=0,$ (2.3)
where the prime means the derivative with respect to the conformal time $\tau$. $\mathcal{H}\equiv a^{\prime}/a$ is the conformal Hubble parameter which can be written as
$\mathcal{H}(\tau)=\frac{1}{\tau}\frac{1+\omega\tau/2}{1+\omega\tau/4},$ (2.4)
where $\omega\equiv\Omega_{m}H_{0}/\sqrt{\Omega_{r}}$ with $H_{0}$ being the Hubble rate today and $\Omega_{m/r}\equiv\bar{\rho}_{m/r,0}/\rho_{c,0}$ is the fractional energy density of matter/radiation today with respect to the present-day critical density $\rho_{c,0}$.
The radiation bath contains photons ($\gamma$), neutrinos ($\nu$), and may also have other light particles in the SM as well as DR. To simplify the discussion, we treat the adiabatic SM radiation bath as one species (denoted as $\gamma_{*}$), which contains all SM radiation components that do not have relative entropy perturbations (defined in Eq. (2.8)). This approach is legitimate because all these components have the same initial conditions and evolution functions. We note that for most cases $\nu$ is a part of $\gamma_{*}$, but we separate it from $\gamma_{*}$ for the NDI case since it contains isocurvature.
Now we move on to study different types of perturbation in cosmology. The gauge invariant curvature perturbation can be written in the Newtonian gauge as [Wands:2000dp]
$\zeta=-\Psi-{\cal H}\frac{\delta\rho}{\rho^{\prime}}~.$ (2.5)
where $\delta\rho=\rho-\bar{\rho}$ denotes the density perturbation. One may also define a curvature fluctuation for each species, namely
$\zeta_{\alpha}=-\Psi-{\cal H}\frac{\delta\rho_{\alpha}}{\rho^{\prime}_{\alpha}}~.$ (2.6)
where $\alpha\in\{\gamma_{*},\nu,b,c,\rm dr\}$ labels the species in the Universe. We further define the perturbation of the total radiation bath as
$\zeta_{r}\equiv\sum_{i\in\,r}R_{i}\zeta_{i},$ (2.7)
where $R_{i}\equiv\bar{\rho}_{i}/\bar{\rho}_{r}$ with $\bar{\rho}_{r}$ being the total background radiation density and $i$ denotes species in the radiation bath. The gauge invariant isocurvature perturbation, or the relative entropy perturbation is defined as
${\cal S}_{\alpha\beta}=3(\zeta_{\alpha}-\zeta_{\beta})~.$ (2.8)
Now we are ready to define initial conditions for all metric and density perturbations in the superhorizon limit $k\tau\ll 1$, where $k$ is the wavenumber of the perturbation. Given the evolution functions from the Einstein equations and Boltzmann equations in the superhorizon limit, there are sets of linearly independent solutions of all perturbations. These solutions are named adiabatic and isocurvature initial conditions. We can then write perturbation variables in each solution as
$X(\mathbf{k},\tau)=c(\mathbf{k})X(k,\tau),$ (2.9)
where $X\in\{\Psi,\Phi,\delta_{\alpha},\theta_{\alpha},\sigma_{\alpha}\}$ with $\delta\equiv\delta\rho/\bar{\rho},\theta,\sigma$ being the density, velocity perturbations, and anisotropic stress respectively. Here $\mathbf{k}\equiv k\hat{\mathbf{k}}$ is the Fourier conjugate to coordinate $\mathbf{x}$. $c(\mathbf{k})$ is a time-independent coefficient which is common for all perturbation variables in a certain set of initial conditions (but different for different initial conditions), which encodes the primordial spacial distribution of perturbations. $X(k,\tau)$ is perturbation-specific but independent of the direction $\mathbf{\hat{k}}$, which is also called the transfer function. For the rest of this section, we show initial conditions for $X(k,\tau)$.
One of the solutions encodes non-vanishing curvature perturbation ($\zeta\neq 0$) with no relative entropy perturbations ($\mathcal{S}_{\alpha\beta}=0$) in the limit $k\tau\to 0$, which is called adiabatic (AD) initial conditions. Fixing $\zeta=1$ and keeping the leading term as $k\tau\to 0$, the adiabatic initial conditions are given by
$\Psi$ $=$ $\Phi=-\frac{2}{3}$~
$\delta_{c}$ $=$ $\delta_{b}=\frac{3}{4}\delta_{\gamma_{*}}=1~~~~~~~~~~~~~~~~~(\textrm{AD})$
$\theta_{c}$ $=$ $\theta_{b}=\theta_{\gamma_{*}}=-\frac{1}{3}k^{2}\tau.$ (2.10)
All other solutions are called isocurvature initial conditions, which have vanishing curvature perturbation $\zeta=0$ but at least one non-zero $\cal S_{\alpha\beta}$. Since there are multiple species in the Universe, we can define different isocurvature initial conditions by different choices of non-vanishing $\cal S_{\alpha\beta}$. In this paper, we choose the adiabatic SM radiation bath $\gamma_{*}$ as the reference and we consider only one non-vanishing $\mathcal{S}_{\alpha\gamma_{*}}$ in each case as $\tau\to 0$. [^2] Since $\mathcal{S}_{\alpha\gamma_{*}}$ is always the same as $\mathcal{S}_{\alpha\gamma}$, it is equivalent to choose the photon bath as the reference. For example, NDI is defined by setting $\mathcal{S}_{\alpha\gamma_{*}}=0$ except for $\mathcal{S}_{\nu\gamma_{*}}\neq 0$. Similarly, we can define BDI ($\mathcal{S}_{b\gamma_{*}}\neq 0$), CDI ($\mathcal{S}_{c\gamma_{*}}\neq 0$), and DRDI ($\mathcal{S}_{\textrm{dr}\gamma_{*}}\neq 0$). We note that the definition of isocurvature conditions is not unique, which depends on the choice of non-vanishing $\cal S_{\alpha\beta}$. Another commonly studied choice is based on $\mathcal{S}_{\alpha r}$ which involves the perturbation of the radiation bath [Bucher:1999re]. We find this choice matches our definition if the isocurvature component is not in the radiation bath like BDI and CDI, but differs for NDI and DRDI. Since $\zeta_{r}$ in the NDI case contains neutrino perturbations that have isocurvature, setting $\mathcal{S}_{br/cr}=0$ does not respect the adiabaticity of baryons and CDM. Therefore, for the study of new physics that generate isocurvature in certain species and the rest remain adiabatic, it is more convenient to use the definition based on $\mathcal{S}_{\alpha\gamma_{*}}$ instead of $\mathcal{S}_{\alpha r}$.
With this definition and setting $\mathcal{S}_{b\gamma_{*}/c\gamma_{*}}=1$, metric and density perturbations in BDI/CDI can be written as (keeping to the order $\omega\tau$)
$\Psi$ $=$ $\Phi=-\frac{1}{8}\omega_{b}\tau$~
$\delta_{c}$ $=$ $\delta_{b}-1=\frac{3}{4}\delta_{\gamma_{*}}=-\frac{3}{8}\omega_{b}\tau~~~~~~~~~~~~~~~(\textrm{BDI})$
$\theta_{c}$ $=$ $\frac{1}{3}\theta_{b}=\frac{1}{3}\theta_{\gamma_{*}}=-\frac{1}{24}\omega_{b}k^{2}\tau^{2}$ (2.11)
$\Psi$ $=$ $\Phi=-\frac{1}{8}\omega_{c}\tau$~
$\delta_{b}$ $=$ $\delta_{c}-1=\frac{3}{4}\delta_{\gamma_{*}}=-\frac{3}{8}\omega_{c}\tau~~~~~~~~~~~~~~~(\textrm{CDI})$
$\theta_{c}$ $=$ $\frac{1}{3}\theta_{b}=\frac{1}{3}\theta_{\gamma_{*}}=-\frac{1}{24}\omega_{c}k^{2}\tau^{2},$ (2.12)
where $\omega_{b/c}\equiv(\Omega_{b/c}/\Omega_{m})\omega$ and $\Omega_{b/c/m}$ is the fractional energy density of baryons/CDM/matter today.
As mentioned before, we separate $\nu$ from $\gamma_{*}$ for NDI. Setting $\mathcal{S}_{\nu\gamma_{*}}=3/4$, [^3] This is the convention commonly used for radiation isocurvature, which ensures the leading isocurvature perturbation (e.g., $\delta_{\nu}$ in NDI) is normalized close to unity [Bucher:1999re]. we get perturbations in NDI as
$\Psi$ $=$ $\Phi=\frac{3}{32}R_{\nu}\omega\tau$~
$\delta_{\gamma_{*}}$ $=$ $\delta_{\nu}-1=\frac{4}{3}\delta_{b/c}=-R_{\nu}+\frac{3}{8}R_{\nu}\omega\tau~~~~~~~~~~~~~~~(\textrm{NDI})$
$\theta_{c}$ $=$ $\frac{1}{3}\theta_{b}=\frac{1}{3}\theta_{\gamma_{*}}=-\frac{1}{24}R_{\nu}\omega k^{2}\tau^{2}.$ (2.13)
Now we are ready to study the interesting relation between NDI and BDI/CDI. To clearly see the relation, we treat the radiation bath as a whole and its perturbation is defined as $\delta_{r}\equiv(1-R_{\nu})\delta_{\gamma_{*}}+R_{\nu}\delta_{\nu}$: $\delta_{r}=\delta_{\gamma_{*}}$ (BDI/CDI) and $\delta_{r}=(3/8)R_{\nu}\omega\tau$ (NDI). Then it is obvious that NDI is simply a rescaling of BDI+CDI:
$X^{\rm NDI}=-\frac{3}{4}R_{\nu}(X^{\rm BDI}+X^{\rm CDI}).$ (2.14)
This can be understood that NDI generates net entropy perturbations between radiation and matter sectors, which effectively act as matter isocurvature with the opposite sign. Since the evolution of BDI and CDI perturbations is almost the same with a simple rescaling, the observational signature of NDI is basically the same as BDI or CDI up to an overall coefficient, which exhibits an interesting “conversion” between NDI and matter isocurvature.
Additional free-streaming DR is introduced for DRDI. Free-streaming DR develops a sizable anisotropic stress, which sources the difference between $\Psi$ and $\Phi$ via the Einstein equations (in radiation domination)
$\Psi-\Phi=6(\mathcal{H}/k)^{2}\sigma,$ (2.15)
where $\sigma$ denotes the total anisotropic stress. Setting $\mathcal{S}_{{\rm dr}\gamma_{*}}=3/4$, perturbations in DRDI are given as (keeping the leading order in $k\tau$)
$\Psi$ $=$ $-\frac{1}{2}\Phi=\frac{(1-R_{\rm dr})R_{\rm dr}}{15+4R_{\rm dr}}$~
$\delta_{\gamma_{*}}$ $=$ $\delta_{\rm dr}-1=\frac{4}{3}\delta_{b/c}=-\frac{R_{\rm dr}(11+8R_{\rm dr})}{15+4R_{\rm dr}}$
$\theta_{c}$ $=$ $-\frac{(1-R_{\rm dr})R_{\rm dr}}{15+4R_{\rm dr}}k^{2}\tau~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(\textrm{DRDI})$
$\theta_{b}$ $=$ $\theta_{\gamma_{*}}=-\frac{19R_{\rm dr}}{60+16R_{\rm dr}}k^{2}\tau$
$\sigma_{\rm dr}$ $=$ $\frac{(1-R_{\rm dr})}{2(15+4R_{\rm dr})}k^{2}\tau^{2}.$ (2.16)
Unlike BDI/CDI or NDI, DRDI has non-vanishing $\Psi$ and $\Phi$ as $\tau\to 0$. This means $\Psi$ and $\Phi$ can be sizable at superhorizon scales, although they are proportional to $R_{\rm dr}$ when $R_{\rm dr}\ll 1$. This leads to qualitatively new features in induced GWs, as we will discuss in the next section.
Figure 1: The absolute value of transfer functions of metric perturbations $|\Phi|$ and $|\Psi|$ in the Newtonian gauge (left) and curvature perturbation $|\zeta|$ (right) as a function of $k\tau$ for AD, CDI and DRDI. $\Phi=\Psi$ for AD and CDI due to no anisotropic stress. $k_{\rm eq}$ is the wavenumber of the mode that enters the horizon at the matter radiation equality, and $R_{\rm dr}\equiv\bar{\rho}_{\rm dr}/\bar{\rho}_{r}$ is the fraction of the DR density to the total radiation density. The scaling of the CDI (DRDI) case is valid for $k\gg k_{\rm eq}$ ($R_{\rm dr}\ll 1$).
We show the time evolution of $\Psi$ and $\Phi$ for different cases in the left panel of figure 1. For illustrative purposes, we just show the AD, CDI and DRDI cases. Since initial conditions for CDI is qualitatively similar to BDI and NDI up to an overall rescaling as mentioned before, these choices covers all relevant features. As seen from figure 1, $\Psi=\Phi\propto C$ for AD and $\Psi=\Phi\propto\omega\tau$ for CDI at superhorizon. They both become damped oscillations after entering the horizon. Around horizon crossing ($k\tau\sim 1$), the CDI case is suppressed by $\omega/k$ (or $k_{\rm eq}/k$ because $k_{\rm eq}=\sqrt{2}\omega$) for large $k$. Therefore we rescale the CDI case by the inverse of this factor. The DRDI case, however, has qualitatively different $\Phi$ and $\Psi$ evolutions due to the presence of anisotropic stress from free-streaming DR. In DRDI, $\Phi,\Psi,\Phi-\Psi\propto R_{\rm dr}$ (see Eq. (2.16)) at superhorizon. They both start damped oscillations after horizon entry and become the same deep inside the horizon due to the anisotropic stress being suppressed (see Eq. (2.15)). Moreover, we present the evolution of the curvature perturbation $\zeta$ in the right panel of figure 1. $\zeta$ of AD starts at one by definition and damps after horizon entry. For isocurvature cases, $\zeta$ starts at zero but grows as time evolves and peaks around horizon entry. We note that for cases without anisotropic stress (AD and CDI), $\zeta$ behaves qualitatively similarly to $\Phi$ or $\Psi$. However, $\zeta$ of DRDI is dramatically different than $\Phi$ or $\Psi$ due to the net anisotropic stress.
Primordial Power Spectrum
Another important information from initial conditions is the initial spatial distribution encoded in $c(\mathbf{k})$ defined in Eq. (2.9). Given the random nature of the initial distribution, we often study observables in terms of this two-point function
$\langle c(\mathbf{k})c(\mathbf{k^{\prime}})\rangle\equiv\frac{2\pi^{2}}{k^{3}}\mathcal{P}(k)(2\pi)^{3}\delta^{3}(\mathbf{k}+\mathbf{k}^{\prime}),$ (2.17)
where $\mathcal{P}(k)$ is the dimensionless power spectrum. For different initial conditions studied in this paper, we get
$\mathcal{P}(k)=\left\{\begin{array}[]{cc}\mathcal{P}_{\zeta}(k)&(\textrm{AD})\\ \mathcal{P}_{\mathcal{S}}(k)&(\textrm{BDI/CDI})\\ \frac{16}{9}\mathcal{P}_{\mathcal{S}}(k)&(\textrm{NDI/DRDI})\end{array}\right$., (2.21)
where $\mathcal{P}_{\zeta}$ and $\mathcal{P}_{\mathcal{S}}$ are the dimensionless power spectra for $\langle\zeta\zeta\rangle$ and $\langle\mathcal{S}\mathcal{S}\rangle$, respectively. Notice that the factor $16/9$ arises because transfer functions for NDI/DRDI modes are defined by setting ${\cal S}=3/4$ as $\tau\to 0$ (see discussions above).
To obtain general constraints on isocurvature from various new physics models, we parameterize $\mathcal{P}(k)$ as the following two general forms: a delta function and a broken power law. The delta function spectrum has the form
$\mathcal{P}_{\rm iso}(k)=A_{\rm iso}\,\delta(\ln k-\ln k_{0}),$ (2.22)
where $A_{\rm iso}$ is the amplitude and $k_{0}$ is the wavenumber where the delta function peaks. While no physical models provide the exact delta function power spectrum, it allows us to isolate the effect of isocurvature on relevant observables at each wavenumber. Therefore, this constraint can be applied to any extended spectrum up to $O(1)$ factors assuming no correlation among different wavenumbers. Another more realistic parameterization is a broken power law spectrum given as
${\cal P}_{\rm iso}(k)=A_{\rm iso}\left\{\begin{matrix}(k/k_{0})^{3}&k\leq k_{0}\\ 1&k>k_{0}\end{matrix}\right$.~, (2.23)
where $k_{0}$ denotes the scale where the transition occurs. The $k^{3}$ dependence for small $k$ is generic for all modes that are outside the horizon around the generation of isocurvature (also known as white noise spectrum, see e.g., [Graham:2015rva, Irsic:2019iff, Feix:2020txt, Amin:2022nlh, Redi:2022llj, Elor:2023xbz, Graham:2024hah, deKruijf:2024voc, Ling:2024qfv, Amin:2025dtd]). If the observation data are mostly from this $k^{3}$ tail (i.e., for large $k_{0}$), the constraint can be applied to any model that has a $k^{3}$ spectrum up to a rescaling. The $k^{0}$ part for $k>k_{0}$ can originate from subhorizon dynamics during inflation, which naturally has nearly scale invariance. One concrete model that has both features is the non-thermal phase transition during inflation [Buckley:2024nen, Barir:2022kzo]. In this case, we can map the parameters in $\mathcal{P}(k)$ to model parameters. The amplitude $A_{\rm iso}\sim\Gamma_{\rm PT}/H_{\rm inf}^{4}$, where $\Gamma_{\rm PT}$ is the phase transition rate per volume and $H_{\rm inf}$ is the Hubble rate during inflation. The comoving scale when this phase transition starts is related to $k_{0}$.
We also show results for the log-normal power spectrum, which is widely studied for induced GWs [Pi:2020otn]. The log-normal spectrum is given as
${\cal P}_{\rm iso}=\frac{A_{\rm iso}}{\sqrt{2\pi}\Delta}\exp\left(-\frac{\ln^{2}(k/k_{0})}{2\Delta^{2}}\right)~,$ (2.24)
where $\Delta$ denotes the dimensionless width. This spectrum can be thought as a smeared version of the delta function spectrum, which can arise in certain inflationary models [Pi:2017gih, Kawasaki:1997ju, Frampton:2010sw, Cai:2018tuh].
Particle Physics Models for Isocurvature Perturbations
To source isocurvature perturbations, it generally needs new sources of perturbations other than the inflaton fluctuations. Additional scalar fields during inflation, such as axion or curvaton models (for reviews, see e.g., [Marsh:2015xka, Mazumdar:2010sa]), can source isocurvature after they decay. Other new dynamics like phase transitions during inflation can also excite isocurvature after they convert the vacuum energy of the phase transition into other species. If the decay or energy transfer is to the decoupled sector, such as CDM and DR, the isocurvature is straightforwardly inherited from the perturbations of the source. However, the BDI and NDI cases are subtler since baryons and neutrinos are tightly coupled to the SM bath and can efficiently transfer energy among them. The key point to maintain net isocurvature in baryons (neutrinos) is that the source must violate baryon (lepton) number, and there are no baryon or lepton number violation interactions in the thermal bath. Therefore the baryon (lepton) number is conserved and maintain the net isocurvature even they are in the thermal equilibrium. Such conditions can be achieved via a late decay of a particle that violates baryon or lepton number after the Electro-Weak sphaleron going out of the bath (when $T\sim 100$ GeV). For example, considering the thermal bath is initially adiabatic with no chemical potential in baryons (leptons), the number violating decays generate a net chemical potential $\mu$ in the phase space distribution of baryons (leptons) and opposite for antibaryons (antileptons). The total energy density of baryons and antibaryons differs from the case without chemical potential (adiabatic case) at $O(\mu^{2})$ for $\mu\ll T$, which becomes the isocurvature. Therefore, in these cases, the amplitude of isocurvature power spectrum is related to the second order of the chemical potential.
Comparing Isocurvature Initial Conditions Before and After Neutrino Decoupling
For comparison, we list the results for conventional definition (with free-steaming neutrinos) of various isocurvature modes in appendix B, but with $\mathcal{S}_{\alpha\gamma_{*}}\neq 0$ instead of $S_{\alpha r}\neq 0$. Here we summarize the main differences between isocurvature modes before and after neutrino decoupling.
  • The main difference comes from the fact that neutrinos are coupled to the SM bath before decoupling, but become free-steaming after decoupling. This leads to vanishing anisotropic stress $\sigma$ for most cases before decoupling, except for DRDI. Therefore, $\Psi=\Phi$ is valid except for DRDI (see Eq. (2.15)). Whereas modes after decoupling always have anisotropic stress from free-steaming neutrinos.
  • For AD, BDI and CDI modes, initial conditions before neutrino decoupling can be simply obtained from ones after neutrino decoupling by setting $R_{\nu}=0$ and removing perturbations from neutrinos.
  • As mentioned before, NDI is qualitatively different before and after neutrino decoupling. Remarkably, NDI is phenomenologically similar to CDI or BDI before decoupling, which reflects an interesting “conversion” between radiation and matter isocurvature.
  • Since free-streaming DR behaves like free-streaming neutrinos, it is equivalent between DRDI before neutrino decoupling and the conventional NDI by exchanging $R_{\rm dr}$ with $R_{\nu}$.
3 Isocurvature Induced Gravitational Waves
With just scalar (metric and density) perturbations, tensor fluctuations $h_{ij}$ in Eq. (2.1) can not be sourced at the first order in perturbation theory due to the mismatch of the spin structure. However, at the second order, there is a source term for tensor modes from the combination of scalar perturbations. Following the standard formalism in Refs. [Kohri2018, Espinosa:2018eve, Pi:2020otn, Gong:2019mui], the equation of motion for tensor perturbation $h_{ij}$ (or gravitational waves) is given by
$h^{\prime\prime}_{ij}+2{\cal H}h^{\prime}_{ij}-\nabla^{2}h_{ij}=S_{ij}~,$ (3.1)
where the source term is [Zhang:2022dgx]
$S_{ij}=$ $4{\cal P}^{ab}{}_{ij}\frac{1}{2}\left[-\partial_{a}\Phi\partial_{b}\Phi+\partial_{a}\Psi\partial_{b}\Phi+\partial_{a}\Phi\partial_{b}\Psi-3\partial_{a}\Psi\partial_{b}\Psi\right$.
$\left.-2\Phi\partial_{a}\partial_{b}\Phi-2\Psi\partial_{a}\partial_{b}\Psi+\frac{4}{3(1+w)}\partial_{a}(\Phi+{\cal H}^{-1}\Psi^{\prime})\partial_{b}(\Phi+{\cal H}^{-1}\Psi^{\prime})\right$.
$\left.-2\partial_{a}\partial_{b}(\Psi-\Phi)\Psi\right]~,$ (3.2)
where ${\cal P}^{ab}{}_{ij}$ is a transverse and traceless operator. For cases with non-vanishing anisotropic stress (e.g., DRDI), there will be additional contributions to the source term $S_{ij}$, which are proportional to the fraction of the density that has anisotropic stress [Weinberg:2003ur, Saga:2014jca, Zhang:2022dgx]. Since we only focus on $R_{\rm dr}\ll 1$ in this study, we neglect these contributions because they are further suppressed by $R_{\rm dr}$.
After the Fourier transform, the equation of motion becomes
$h^{\prime\prime}_{\mathbf{k},\lambda}+2{\cal H}h^{\prime}_{\mathbf{k},\lambda}+k^{2}h_{\mathbf{k},\lambda}=S_{\mathbf{k},\lambda}$~ (3.3)
with $\lambda$ denoting two polarizations of $h_{ij}$ and
$S_{\mathbf{k},\lambda}=4\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3}}e^{ij}_{\mathbf{k},\lambda}q_{i}q_{j}{c(\mathbf{q})}c(|\mathbf{k}-\mathbf{q}|)f(\tau,q,|\mathbf{k}-\mathbf{q}|)~,$ (3.4)
where $e^{ij}_{\mathbf{k},\lambda}$ is the polarization tensor and
$f(\tau,q,|\mathbf{k}-\mathbf{q}|)$ $=$ $\frac{1}{2}\Phi(q,\tau)\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{1}{2}\Psi(q,\tau)\Psi(|\mathbf{k}-\mathbf{q}|,\tau)$ (3.5)
$+$ $\frac{1}{2}\left(\Phi(q,\tau)+\frac{\Psi^{\prime}(q,\tau)}{\cal H}\right)\left(\Phi(|\mathbf{k}-\mathbf{q}|,\tau)+\frac{\Psi^{\prime}(|\mathbf{k}-\mathbf{q}|,\tau)}{\cal H}\right)~.$
Here $c$ and $\Psi,\Phi$ are defined in Eq. (2.9) and initial conditions for $\Psi,\Phi$ are given in section 2.
The energy density of gravitational waves relevant for observations is given by [Domenech:SIGWRe]
$\rho_{\rm GW}=\frac{M_{\rm pl}^{2}}{4a^{2}}\langle h^{\prime ij}h^{\prime}_{ij}\rangle,$ (3.6)
where $M_{\rm pl}$ is the Planck mass. This definition is valid when all relevant tensor modes are well inside the horizon, and we have used the approximation $h^{\prime}\approx kh$. The two-point function in the Fourier space can be written as
$\langle h_{\mathbf{k},\lambda}h_{\mathbf{k}^{\prime},\lambda}\rangle\equiv\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{h,\lambda}(k)(2\pi)^{3}\delta^{3}(\mathbf{k}+\mathbf{k}^{\prime}),$ (3.7)
where $\mathcal{P}_{h}(k)$ is the dimensionless power spectrum of $h$. It is common to use the spectral density fraction ($\Omega_{\rm GW}$) defined as
$\Omega_{\rm GW}(k)\equiv\frac{1}{\bar{\rho}}\frac{d\rho_{\textrm{GW}}}{d\ln k}=\frac{k^{2}}{12\mathcal{H}^{2}}\overline{\mathcal{P}}_{h}(k),$ (3.8)
where $\overline{\cal P}_{h}\equiv\sum_{\lambda}\overline{\mathcal{P}}_{h,\lambda}$ and the overline denotes the average over oscillations (see below).
Figure 2: The $f(x,u,v)$ function in the source term for induced GWs for AD (blue), CDI (orange) and DRDI (green) with $x=k\tau$. For simplicity we chose $u=v=1$ but the main behavior is independent of such particular values.
The key to obtain GW density is to calculate $\overline{\mathcal{P}}_{h}$. Using the Green’s function method and assuming no primordial sources of GWs, we can write the solution of $h_{\mathbf{k},\lambda}$ as
[…]
where $G(\tau,\tilde{\tau})=k\tilde{\tau}^{2}[y_{0}(k\tau)j_{0}(k\tilde{\tau})-j_{0}(k\tau)y_{0}(k\tilde{\tau})]=\sin(k(\tau-\tilde{\tau}))\times\tilde{\tau}/(k\tau)$ is the Green’s function of Eq. (3.3) in the radiation dominated universe with $j_{0}$ and $y_{0}$ being the spherical Bessel functions. We can then write the expression for $\overline{\mathcal{P}}_{h}$ as
$\overline{\cal P}_{h}=8\int_{0}^{\infty}dv\int_{|1-v|}^{1+v}du\left(\frac{4v^{2}-(1-u^{2}+v^{2})^{2}}{4uv}\right)^{2}\overline{I^{2}(x,v,u)}{\cal P}(ku){\cal P}(kv)~,$ (3.10)
where we have changed variables $x=k\tau,q=vk,|\mathbf{k}-\mathbf{q}|=uk$. $\mathcal{P}(k)$ is the primordial power spectrum defined in Eq. (2.17). To get Eq. (3.10), we assume the primordial perturbations are mostly Gaussian, and thus the four-point function of $c$ can be rewritten as the product of two-point functions.
Furthermore, we have inserted all the time dependence into a kernel function defined by
$I(x,u,v)=k\int_{0}^{x}d\tilde{x}\ G(x,\tilde{x})f(\tilde{x},u,v)~,$ (3.11)
where $f(x,u,v)$ is given in Eq. (3.5). Plugging in the Green’s function of gravitational waves, the kernel function can be further split as
$I(x\gg 1,u,v)=I_{y}(u,v)y_{0}(x)+I_{j}(u,v)j_{0}(x)~,$ (3.12)
where
$I_{y/j}=\int_{0}^{\infty}d\tilde{x}\ \tilde{x}^{2}\left\{\begin{matrix}j_{0}(\tilde{x})\\ y_{0}(\tilde{x})\end{matrix}\right\}f(\tilde{x},u,v)~.$ (3.13)
Since we are interested in the results deep inside the horizon $x=k\tau\gg 1$, $I(x,u,v)$ is a highly oscillatory function due to $y_{0}(x)$ and $j_{0}(x)$. To get a meaningful result, we take a time average over oscillations around $x$ and obtain $\overline{y^{2}_{0}(x)}=\overline{j^{2}_{0}(x)}=1/(2x^{2})$, $\overline{y_{0}(x)j_{0}(x)}=0$. This also reflects the fact that the measurement of stochastic GW background can not resolve the real time evolution but just the time-averaged power. Therefore, we can get
$\overline{I^{2}(x,u,v)}=\frac{I_{y}^{2}(u,v)+I_{j}^{2}(u,v)}{2x^{2}}.$ (3.14)
Combining Eqs. (3.8), (3.10) and (3.14) as well as $\mathcal{H}\approx\tau^{-1}$ in radiation domination, we can infer that $\Omega_{\rm GW}(k)$ is independent of $\tau$ around the generation. Taking into account the subsequent heating of the radiation bath from changing degrees of freedom, the observed GW density today is
$h^{2}\Omega_{\rm GW,0}(k)=h^{2}\Omega_{r,0}\left(\frac{g_{*}(T_{\rm GW})}{g_{*}(T_{0})}\right)\left(\frac{g_{*S}(T_{0})}{g_{*S}(T_{\rm GW})}\right)^{4/3}\Omega_{\rm GW}(k)=3.0\times 10^{-5}\left(\frac{17.25}{g_{*}(T_{\rm GW})}\right)^{1/3}\Omega_{\rm GW}(k),$
where $h\equiv H_{0}/(100\,\textrm{km/s/Mpc})$, $g_{*}(T)$ and $g_{*S}(T)$ are the effective number of relativistic degrees of freedom for energy and entropy, respectively. $T_{0}$ is the present-day temperature and $T_{\rm GW}$ is the temperature when the GWs are generated, which is $T_{\rm GW}=O(100)$MeV in this study. Therefore, $g_{*}(T_{\rm GW})=g_{*S}(T_{\rm GW})=17.25$, which includes the contribution from pions, muons, and electrons/positions. The radiation density today $\Omega_{r,0}h^{2}=4.2\times 10^{-5}$ is taken from [Planck:2018vyg].
To explicitly view the effects on induced GWs from different initial conditions, we should look at the source term, in particular $f(x,u,v)$ defined in Eq. (3.5) which encodes all the time evolution. As can be seen from figure 2, the subhorizon dynamics are rather universal, which exhibit damped oscillations. However, the superhorizon behavior depends on the certain initial condition. Both the adiabatic and DRDI cases remain constant at superhorizon $x\ll 1$, whereas the CDI case exhibits a $\tau^{2}$ dependence. These features originate from the fact that $\Phi$ and $\Psi$ are constant for AD and DRDI but $\propto\tau$ for CDI outside the horizon. We note that the CDI case is further suppressed by $(k_{\rm eq}/k)^{2}$ and the DRDI case by $R_{\rm dr}^{2}$, which reflects the suppression of the density of the corresponding species. Therefore, for modes relevant for PTA observations ($k\gg k_{\rm eq}$), the source term of CDI is highly suppressed but that of DRDI can still be sizable for appropriate $R_{\rm dr}^{2}$. The cases of BDI and NDI are similar to CDI with an overall rescaling, and thus we don’t show them here.
Given the isocurvature power spectrum $\mathcal{P}(k)$ in Eqs. (2.22), (2.23) and (2.24), we can obtain the spectrum of induced GWs ($\Omega_{\rm GW}$). As seen from figure 3, the results for AD and DRDI are almost the same for a given power spectrum, up to an overall coefficient $\sim R_{\rm dr}^{4}$ for DRDI. This result is consistent with the behavior of $f(x,u,v)$ discussed before. However, there are subtle differences for the delta-function power spectrum. There is a dip to the left side of the peak of the GW spectrum for AD case, but not for DRDI. This feature has been smeared out for the smooth $\mathcal{P}(k)$. The CDI case has a very different shape for large $k$, which is the result of suppression of matter density at early times. For the broken power law $\mathcal{P}(k)$ in the flat region ($k>k_{0}$), both AD and DRDI cases stay constant but CDI $\propto k^{-4}$. Therefore, the GW spectrum induced by DRDI is dramatically different from the previously studied CDI case. This provides the opportunities to study new GW spectra from isocurvature induced GWs, and the constraints on new physics with the existing GW data.
Figure 3: The spectrum of GWs ($\Omega_{\rm GW}$) induced by AD (blue), CDI (orange) and DRDI (green) perturbations. We consider three choices of power spectrum: a delta function (Eq. (2.22)), a log-normal form with $\Delta=0.1$ (Eq. (2.24)) and a broken power law (Eq. (2.23)). For each power spectrum, $k_{0}$ denotes the characteristic scale and we have set the amplitude $A_{\rm ad/iso}=1$. Here we choose $k_{0}\gg k_{\rm eq}$ for CDI and $R_{\rm dr}\ll 1$ for DRDI.
Curvature vs Isocurvature Induced GWs
Here we would like to clarify some intuitions for curvature and isocurvature induced GWs. One common intuition of SIGWs is that GWs are always sourced by curvature perturbations. In the case of isocurvature initial conditions, the curvature perturbation $\zeta$ starts at zero (in the limit $\tau\to 0$) by definition, but grows and peaks around the horizon crossing. Therefore isocurvature induced GWs are mostly generated around the horizon crossing where the curvature perturbation peaks and the density of induced GWs is also determined by the size of curvature perturbations: $\Omega_{\rm GW}\propto\zeta^{4}|_{k\tau\sim 1}$.
We argue that this intuition is only parametrically correct, but may obscure some interesting features in certain cases. We emphasize that it is the source term in Eq. (3.4) that generates GWs, which is a non-trivial function involving the combination of $\Psi$ and $\Phi$, and their derivatives. Therefore, it is not just curvature perturbations as the source. For the cases without anisotropic stress (including AD, BDI, CDI and NDI before neutrino decoupling), $\Phi=\Psi$ and they have a similar time dependence as $\zeta$ (see figure 1). Therefore, the source term can be roughly rewritten in terms of the curvature perturbation $\zeta$. This explains the suppression of $\Omega_{\rm GW}$ in the BDI/CDI/NDI case compared to the AD case. The curvature perturbation $\zeta\propto\omega/k$ in BDI/CDI/NDI around the horizon crossing $k\tau\sim 1$, which is parametrically suppressed compared to the the AD case ($\zeta=O(1)$) for large $k$. This suppression can also be understood from the fact that the matter density is suppressed at early times. Given that GWs are generated mainly around the horizon crossing, $\Omega_{\rm GW}$ in the BDI/CDI/NDI case is suppressed by $(\omega/k)^{4}\propto\zeta^{4}|_{k\tau\sim 1}$. Therefore, it matches the common intuition that isocurvature induced GWs are suppressed because the curvature perturbation in these cases are suppressed.
However, this intuition does not capture all interesting features of DRDI induced GWs. Due to the net anisotropic stress from free-streaming DR, the superhorizon evolution of metric perturbations ($\Phi$ and $\Psi$) is dramatically different from that of $\zeta$ (see figure 1). Therefore, unlike the previous cases, the source term cannot be simply rewritten in terms of $\zeta$. For example, the source term is not suppressed at $\tau\to 0$ even though $\zeta\to 0$ (see figure 2). The induced GWs from DRDI are qualitatively different from other isocurvature cases due to the presence of anisotropic stress, and the density of induced GWs can be much larger than those from CDI since DR density ratio ($R_{\rm dr}$) stay fixed in the radiation domination. Although the usual intuition $\Omega_{\rm GW}\propto\zeta^{4}|_{k\tau\sim 1}\propto R_{\rm dr}^{4}$ (for $R_{\rm dr}\ll 1$) still parametrically holds, it totally obscures the superhorizon dynamics. Therefore, one needs to study the full source term to understand induced GWs for general isocurvature cases.
4 Constraints from Pulsar Timing Arrays
Figure 4: The best fit and constraint of AD power spectrum from NANOGrav. We utilize two general parametrizations of the isocurvature power spectrum: a delta function (left, Eq. (2.22)) and a broken power law (right, Eq. (2.23)). Each spectrum has a characteristic scale $k_{0}$ and amplitude $A_{\rm ad}$. The constraint is shown in the gray shaded region, while $1\sigma,\ 2\sigma$ CL regions are depicted in progressively lighter shades of blue.
Having studied the GW spectrum induced by primordial adiabatic and various isocurvature perturbations, we will perform data fitting with the NANOGrav 15-year dataset [NANOGrav:2023gor, NANOGrav:2023hde, NANOGrav:2023hvm, nano_grav_2023_dataset]. In this section, we will present best fit and constraints for AD, CDI and DRDI assuming only one mode dominates for each case. Since BDI and NDI have similar observational signals as CDI, their results can be simply recast from CDI with an overall rescaling. For our data analysis, we employ the parameter $k_{0}$ and the power spectrum amplitude $A_{\rm ad/iso}$ for each mode. We apply the Bayesian inference method to determine the best fit of SIGWs. We adopt PTArcade [andrea_mitridate_2023, Mitridate:2023oar] to sample the posterior probability and generate upper limits above which the additional model is “strongly disfavored” according to the Jeffreys scale [NANOGrav:2023hvm, Jeffreys]. The priors of parameters follow log-uniform distributions within $\log_{10}(k_{0}/{\rm Mpc}^{-1})\in[4,10]$, $\log_{10}(A_{\rm ad})\in[-3,1]$, $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})\in[-3,1]$, and $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})\in[-3,1]$.
We first show the constraints and best fit for adiabatic (curvature) perturbations (see figure 4). Constraints on curvature perturbations from the latest PTA data have been discussed in several recent studies [NANOGrav:2023hvm, Ellis:2023oxs, Cai:2023dls, Figueroa:2023zhu, Yi:2023tdk, Yi:2023mbm, You:2023rmn, Franciolini:2023pbf, Liu:2023ymk, Liu:2023pau, Liu:2023hpw, Wang:2023ost, Kumar:2024hsi, Chen:2024fir]. Our results with a delta function spectrum agree with those existing results, which serves as the validation of our data analysis. We then present new results for adiabatic perturbations with the broken power law spectrum (see Eq. (2.23)). Current data indicate that the amplitude of the adiabatic power spectrum $A_{\rm ad}\lesssim 10^{-2}$ around $k_{0}\sim 10^{6}\,\textrm{Mpc}^{-1}$ for the delta function power spectrum, or $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ for the broken power law spectrum.
The results for CDI and DRDI with the delta function power spectrum exhibit similar features as the AD case. They all have the strongest constraining power around $k_{0}\sim 10^{6}\,\textrm{Mpc}^{-1}$, but the sensitivity drops sharply for $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ and gradually decreases for $k_{0}\gtrsim 10^{6}\,\textrm{Mpc}^{-1}$. This reflects the shape of induced GW spectrum (figure 3), which has a peak around $k_{0}$, a sharp edge for large $k$ and a long tail for small $k$. The differences are the peak constraint reads $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})\lesssim-2.05$ for CDI and $\log_{10}(R_{\rm dr}^{2}A_{\rm iso})\lesssim-1.57$ for DRDI.
The constraint on DRDI with a broken power law spectrum is flat for $k_{0}\lesssim 10^{6}\,\textrm{Mpc}^{-1}$ and weakens for large $k_{0}$ (see figure 5). These features originate from the DRDI GW spectrum (figure 3) which has a plateau for $k\gtrsim k_{0}$ and decays for $k\lesssim k_{0}$. However, the CDI constraint on $\log_{10}((k_{\rm eq}/k_{0})^{2}A_{\rm iso})$ has a different low-$k$ behavior that scales as $k_{0}^{-2}$. This is because the high-$k$ region of the GW spectrum of CDI scales as $k^{-4}$.
Figure 5: The best fit and constraint of DRDI and CDI power spectrum from NANOGrav. We utilize two general parametrizations of the isocurvature power spectrum: a delta function (left, Eq. (2.22)) and a broken power law (right, Eq. (2.23)). Each spectrum has a characteristic scale $k_{0}$ and amplitude $A_{\rm iso}$. The constraint is shown in the gray shaded region, while $1\sigma,\ 2\sigma$ CL regions are depicted in progressively lighter shades of blue.
The best fit for all cases with delta function power spectrum has a peak at $k_{0}\gg 10^{6}\,\textrm{Mpc}^{-1}$, which indicates that the NANOGrav data only probe the low-$k$ tail of the induced GW spectrum. We show the best fit of the DRDI case as an example in figure 6. As clearly seen in this figure, the peak of the best fit of the delta function power spectrum lies beyond the range of the plot. Interestingly, the posterior of the broken power law case has a bimodal feature. Both AD and DRDI cases have similar features but DRDI is more pronounced. We show the spectrum of GWs from two peaks of the posterior of DRDI in figure 6 with BPL1 ($\log_{10}(k_{0}/\textrm{Mpc}^{-1})=6.82,\log_{10}(R_{\rm dr}^{2}A_{\rm iso})=-1.00$) and BPL2 ($\log_{10}(k_{0}/\textrm{Mpc}^{-1})=7.46,\log_{10}(R_{\rm dr}^{2}A_{\rm iso})=-0.48$). This bimodal feature originates from the fact that the induced GW spectrum for DRDI exhibits “two steps” around the transition scale $k_{0}$ (see figure 3). In particular, the BPL1 with a lower $k_{0}$ fits the data around $f\sim 10^{8}\rm Hz$ well, but not for lower frequencies. In contrast, the BPL2 fits the low frequency data better but misses the features of the data around $f\sim 10^{8}\rm Hz$. It will be interesting to study future high precision data to distinguish these two models. Note that this bimodal feature originates from the sharp transition in the broken power law spectrum at $k_{0}$. A smoother transition will obscure this feature by smearing the induced GW spectrum.
We further compare the constraints on the isocurvature power spectrum from induced GWs with other observations such as the CMB+BAO, Lyman-$\alpha$ forest and CMB spectral distortions. We can see from figure 7 that our constraint from NANOGrav 15 year data places a stringent constraint around $k_{0}\sim 10^{6}\textrm{Mpc}^{-1}$, which are complementary to other constraints on larger scales. We acknowledge that a few other strong constraints on isocurvature are not present in this figure because they are derived assuming either BDI or CDI (see e.g., [Passaglia:2021jla, Graham:2024hah, Inomata:2018htm, Bagherian:2025puf]).
Figure 6: The spectrum of DRDI induced GWs for the best fit parameters with NANOGrav 15 year data. The result for the delta function power spectrum is shown in green. Since the case with the broken power law spectrum has a bimodal feature, we show the results for these two peaks and label them as BPL1 (blue) and BPL2 (orange). The data with error bars from NANOGrav are shown in gray.
Finally, we comment on two features of the constraints on isocurvature. Firstly, the combination $R_{\rm dr}^{2}A_{\rm iso}$ for DRDI and $(k_{\rm eq}/k_{0})^{2}A_{\rm iso}$ for CDI controls the physical effect of the perturbation. This means the break-down of the perturbation theory is not $A_{\rm iso}\gtrsim 1$, but $R_{\rm dr}^{2}A_{\rm iso}\gtrsim 1$ or $(k_{\rm eq}/k_{0})^{2}A_{\rm iso}\gtrsim 1$. For most regions of the constraints derived in this study is safe from this consideration. However, $A_{\rm iso}\gg 1$ means that the isocurvature distribution must be non-Gaussian. The reason is the following. The density contrast $\delta_{\alpha}\geq-1$ due to $\delta\rho_{\alpha}+\bar{\rho}_{\alpha}\geq 0$. If the isocurvature amplitude $A_{\rm iso}\sim\langle\delta_{\alpha}\delta_{\alpha}\rangle\gg 1$, the distribution of $\delta_{\alpha}$ must be mostly around 1-1 to allow for substantial large positive values. Therefore, constraints from our study assuming Gaussian isocurvature fluctuations may need to be modified by non-Gaussian effects given a concrete model.
Figure 7: Constraints on isocurvature power spectrum for DRDI, CDI and BDI cases with two general parametrizations: a delta function (left, Eq. (2.22)) and a broken power law (right, Eq. (2.23)). Our limit from NANOGrav 15 year data is shown in blue. Gray shaded regions are excluded by the CMB, BAO, Lyman-$\alpha$ forest and CMB spectral distortions [Buckley:2025zgh]. The dotted line indicates the projected sensitivity for the future experiment PIXIE probing CMB distortions [Kogut:2011xw].
5 Conclusions
[…]
Acknowledgement
We thank Lang Liu and Soubhik Kumar for useful discussions on this work and feedback on a draft of this manuscript. This work was supported in part by the National Key R&D Program of China (2021YFC2203100), by the National Natural Science Foundation of China (12433002, 12261131497), by CAS young interdisciplinary innovation team (JCTD-2022-20), by 111 Project (B23042), by CSC Innovation Talent Funds, by USTC Fellowship for International Cooperation, and by USTC Research Funds of the Double First-Class Initiative. PD is supported by the National Natural Science Foundation of China (Grants No. T2388102). JZ is supported by the National Natural Science Foundation of China (125B1023).
Appendix A Evolution Functions for Primordial Perturbations
In this appendix, we show the set of Einstein equations and Boltzmann equations that govern the evolution of primordial perturbations. In the Newtonian gauge with the notation defined in section 2, the Einstein equations can be written as [Kodama:1986fg, Kodama:1986ud, Ma:1995ey]
$\mathcal{H}^{-1}\Psi^{\prime}+\Phi+\frac{1}{3}\left(\frac{k}{\cal H}\right)^{2}\Psi$ $=$ $-\frac{1}{2}\delta$ (A.1)
$\frac{k^{2}}{\mathcal{H}^{2}}\Psi^{\prime}+\frac{k^{2}}{\mathcal{H}}\Phi$ $=$ $\frac{3}{2}(1+w)\theta$ (A.2)
$\mathcal{H}^{-2}\Psi^{\prime\prime}+\mathcal{H}^{-1}\left(\Phi^{\prime}+2\Psi^{\prime}\right)+\left(1+\frac{2{\cal H}^{\prime}}{{\cal H}^{2}}\right)\Phi-\frac{1}{3}\left(\frac{k}{\cal H}\right)^{2}(\Phi-\Psi)$ $=$ $\frac{3}{2}\frac{\delta p}{\bar{\rho}}$ (A.3)
$\left(\frac{k}{\mathcal{H}}\right)^{2}(\Psi-\Phi)$ $=$ $\frac{9}{2}(1+w)\sigma,$ (A.4)
where $w\equiv\bar{p}/\bar{\rho}$ and $\delta p\equiv p-\bar{p}$. We can also relate the total perturbations to the weighted sum of individual ones: $\delta=(\sum_{\alpha}\bar{\rho}_{\alpha}\delta_{\alpha})/\bar{\rho}$, $\theta=(\sum_{\alpha}(\bar{\rho}_{\alpha}+\bar{p}_{\alpha})\theta_{\alpha})/(\bar{\rho}+\bar{p})$, $\sigma=(\sum_{\alpha}(\bar{\rho}_{\alpha}+\bar{p}_{\alpha})\sigma_{\alpha})/(\bar{\rho}+\bar{p})$. For the cases with no anisotropic stress (AD/BDI/CDI/NDI), we have $\Phi=\Psi$ from the last equation of Eq. (A.4). Therefore, we only need to solve a single equation to get the evolution for $\Phi$ and $\Psi$.
To obtain the Boltzmann equations for density perturbations, we consider baryons are strongly coupled with the SM radiation bath (denoted as $\gamma_{*}$) and use the tight-coupling approximation: setting $\theta_{b}=\theta_{\gamma_{*}}$ and neglecting the anisotropic stress $\sigma_{\gamma_{*}}=0$ [Ma:1995ey, Bucher:1999re]. With this approximation, the evolution functions of $\theta_{b}$ and $\theta_{\gamma_{*}}$ can be combined into one equation. In the radiation dominated era ($w=1/3$), the Boltzmann equations for the density perturbations of different species read
$\delta_{\gamma_{*}}^{\prime}$ $=$ $-\frac{4}{3}\theta_{\gamma_{*}}+4\Psi^{\prime}~,$ (A.5)
$\delta_{c}^{\prime}$ $=$ $-\theta_{c}+3\Psi^{\prime}~,$ (A.6)
$\delta_{b}^{\prime}$ $=$ $-\theta_{\gamma_{*}}+3\Psi^{\prime}~,$ (A.7)
$(1+r_{b})\theta_{\gamma_{*}}^{\prime}$ $=$ $\frac{1}{4}k^{2}\delta_{\gamma_{*}}-r_{b}{\cal H}\theta_{\gamma_{*}}+(1+r_{b})k^{2}\Phi~,$ (A.8)
$\theta_{c}^{\prime}$ $=$ $-\mathcal{H}\theta_{\rm c}+k\Phi~,$ (A.9)
where $r_{b}\equiv 3\bar{\rho}_{b}/(4\bar{\rho}_{\gamma_{*}})$. For NDI, we separate $\nu$ from $\gamma_{*}$ (as mentioned in section 2) but $\nu$ has the same evolution functions as $\gamma_{*}$, so we don’t show it explicitly here.
For DRDI with free-streaming dark radiation, we need to add an additional DR species which contains net anisotropic stress. The complete Boltzmann hierarchy for DR involves an arbitrarily large $\ell$ moment (typically denoted as $F_{\ell}$):
$\delta_{\rm dr}^{\prime}$ $=$ $-\frac{4}{3}\theta_{\rm dr}+4\Psi^{\prime}~,$ (A.10)
$\theta_{\rm dr}^{\prime}$ $=$ $\frac{1}{4}k^{2}\delta_{\rm dr}-k^{2}\sigma_{\rm dr}+k^{2}\Phi~,$ (A.11)
$F_{\rm dr,\ell}^{\prime}$ $=$ $\frac{k}{2\ell+1}\left[\ell F_{\rm dr,\ell-1}-(\ell+1)F_{\rm dr,\ell+1}\right]~~~~(\ell\geq 2),$ (A.12)
where $\delta=F_{0}$, $\theta=(3/4)kF_{1}$, $\sigma=(1/2)F_{2}$. The effects of high $\ell$ moments on observables are suppressed by powers of $k\tau$ in the superhorizon limit. The Boltzmann hierarchy can therefore be simplified by truncating it at a certain $\ell_{\rm max}$ depending on the desired precision. In this study, we choose $\ell_{\rm max}=8$ and obtain closed differential equations using $F_{\ell_{\rm max}+1}\approx(2\ell_{\rm max}+1)F_{\ell_{\rm max}}/(k\tau)-F_{\ell_{\rm max}-1}$ [Ma:1995ey] and initial conditions $F_{\ell\geq 3}=0$.
Appendix B Isocurvature Initial Conditions After Neutrino Decoupling
Here we show isocurvature initial conditions after neutrino decoupling, where neutrinos are treated as free-streaming radiation. These cases are widely studied in the literature for the physics relevant for the CMB and large scale structures. The conventional definition of isocurvature is based on setting $\zeta=0$ and a certain $\mathcal{S}_{\alpha r}\neq 0$ [Bucher:1999re]. As argued in section 2, we choose $\mathcal{S}_{\alpha\gamma}\neq 0$ instead of $\mathcal{S}_{\alpha r}\neq 0$ to fix isocurvature initial conditions. With this definition, CDI and BDI are the same, but NDI and DRDI are slightly different than those in the literature. We also list the adiabatic initial conditions for comparison.
In the Newtonian gauge and keeping the leading order in $\tau$, the adiabatic initial conditions are given by
$\Psi$ $=$ $-\frac{10+R_{\nu}}{15+4R_{\nu}}~~,~~\Phi=-\frac{10}{15+4R_{\nu}}$~
$\delta_{c}$ $=$ $\delta_{b}=\frac{3}{4}\delta_{\gamma}=\frac{3}{4}\delta_{\nu}=\frac{15}{15+4R_{\nu}}~~~~~~~~~~~~~~~(\textrm{AD})$
$\theta_{c}$ $=$ $\theta_{b}=\theta_{\gamma}=\theta_{\nu}=-\frac{5}{15+4R_{\nu}}k^{2}\tau$
$\sigma_{\nu}$ $=$ $\frac{2}{3(15+4R_{\nu})}k^{2}\tau^{2}~,$ (B.1)
where $R_{\nu}\equiv\bar{\rho}_{\nu}/\bar{\rho}_{r}$.
The BDI and CDI initial conditions are
$\Psi$ $=$ $-\frac{1}{8}\frac{15+4R_{\nu}}{15+2R_{\nu}}\omega_{b}\tau~,~\Phi=-\frac{1}{8}\frac{15-4R_{\nu}}{15+2R_{\nu}}\omega_{b}\tau$~
$\delta_{c}$ $=$ $\delta_{b}-1=\frac{3}{4}\delta_{\gamma}=\frac{3}{4}\delta_{\gamma_{\nu}}=-\frac{3}{8}\frac{15+4R_{\nu}}{15+2R_{\nu}}\omega_{b}\tau~~~~~~~~~~~~~~~(\textrm{BDI})$
$\theta_{b}$ $=$ $\theta_{\gamma}=\theta_{\nu}=-\frac{1}{8}\frac{15}{15+2R_{\nu}}\omega_{b}k^{2}\tau^{2}~,~\theta_{c}=-\frac{1}{24}\frac{15-4R_{\nu}}{15+2R_{\nu}}\omega_{b}k^{2}\tau^{2}$
$\sigma_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{b}k^{2}\tau^{3}~,$ (B.2)
$\Psi$ $=$ $-\frac{1}{8}\frac{15+4R_{\nu}}{15+2R_{\nu}}\omega_{c}\tau~,~\Phi=-\frac{1}{8}\frac{15-4R_{\nu}}{15+2R_{\nu}}\omega_{c}\tau$~
$\delta_{b}$ $=$ $\delta_{c}-1=\frac{3}{4}\delta_{\gamma}=\frac{3}{4}\delta_{\gamma_{\nu}}=-\frac{3}{8}\frac{15+4R_{\nu}}{15+2R_{\nu}}\omega_{c}\tau~~~~~~~~~~~~~~~(\textrm{CDI})$
$\theta_{b}$ $=$ $\theta_{\gamma}=\theta_{\nu}=-\frac{1}{8}\frac{15}{15+2R_{\nu}}\omega_{c}k^{2}\tau^{2}~,~\theta_{c}=-\frac{1}{24}\frac{15-4R_{\nu}}{15+2R_{\nu}}\omega_{c}k^{2}\tau^{2}$
$\sigma_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{c}k^{2}\tau^{3}$ (B.3)
where $\omega_{b/c}\equiv\Omega_{b/c}H_{0}/\sqrt{\Omega_{r}}$ and $\Omega_{b/c/r}$ is the fractional energy density of baryons/CDM/radiation today.
The NDI is given by
$\Psi$ $=$ $\frac{R_{\nu}(1-R_{\nu})}{15+4R_{\nu}}~,~\Phi=\frac{-2R_{\nu}(1-R_{\nu})}{15+4R_{\nu}}$~
$\delta_{\gamma}$ $=$ $\frac{4}{3}\delta_{c}=\frac{4}{3}\delta_{b}=-R_{\nu}\frac{11+8R_{\nu}}{15+4R_{\nu}}$
$\delta_{\nu}$ $=$ $\frac{(15+8R_{\nu})(1-R_{\nu})}{15+4R_{\nu}}~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(\textrm{NDI})$
$\theta_{b}$ $=$ $\theta_{\gamma}=-\frac{19R_{\nu}}{4(15+4R_{\nu})}k^{2}\tau$~
$\theta_{\nu}$ $=$ $\frac{15(1-R_{\nu})}{4(15+4R_{\nu})}k^{2}\tau~,~\theta_{c}=-\frac{R_{\nu}(1-R_{\nu})}{15+4R_{\nu}}k^{2}\tau$~
$\sigma_{\nu}$ $=$ $\frac{(1-R_{\nu})}{2(15+4R_{\nu})}k^{2}\tau^{2}~.$ (B.4)
The DRDI is given by
$\Psi$ $=$ $-\frac{1}{2}\Phi=\frac{R_{\rm dr}(1-R_{\rm dr}-R_{\nu})}{15+4R_{\rm dr}+4R_{\nu}}$~
$\delta_{\gamma}$ $=$ $\delta_{\nu}=\frac{4}{3}\delta_{b}=\frac{4}{3}\delta_{c}=-\frac{R_{\rm dr}(11+8R_{\rm dr}+8R_{\nu})}{15+4R_{\rm dr}+R_{\nu}}$
$\delta_{\rm dr}$ $=$ $\frac{15-7R_{\rm dr}-8R_{\rm dr}^{2}+4R_{\nu}-8R_{\rm dr}R_{\nu}}{15+4R_{\rm dr}+4R_{\nu}}$
$\theta_{b}$ $=$ $\theta_{\gamma}=\theta_{\nu}=-\frac{19R_{\rm dr}}{4(15+4R_{\rm dr}+4R_{\nu})}k^{2}\tau~~~~~~~~~~~~~~~~~~~~(\textrm{DRDI})$
$\theta_{c}$ $=$ $-\frac{R_{\rm dr}(1-R_{\rm dr}-R_{\nu})}{15++4R_{\rm dr}+4R_{\nu}}k^{2}\tau~,~\theta_{\rm dr}=\frac{15-15R_{\rm dr}+4R_{\nu}}{4(15+4R_{\rm dr}+4R_{\nu})}k^{2}\tau$
$\sigma_{\nu}$ $=$ $-\frac{19R_{\rm dr}}{30(15+4R_{\rm dr}+4R_{\nu})}k^{2}\tau^{2}~,~\sigma_{\rm dr}=\frac{15-15R_{\rm dr}+4R_{\nu}}{30(15+4R_{\rm dr}+4R_{\nu})}k^{2}\tau^{2}~.$ (B.5)
We also show these initial conditions in the synchronous gauge. The perturbed FRW metric in the synchronous gauge is written as
$ds^{2}=a(\tau)^{2}\left[-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\right].$ (B.6)
The metric perturbation $h_{ij}$ can be written in Fourier space as
$h_{ij}(\mathbf{k},\tau)=\left[\hat{k}_{i}\hat{k}_{j}h(\mathbf{k},\tau)+\left(\hat{k}_{i}\hat{k}_{j}-\frac{1}{3}\delta_{ij}\right)6\eta(\mathbf{k},\tau)\right],$ (B.7)
where $h$ and $\eta$ denote the trace and traceless longitudinal part of $h_{ij}$ respectively.
In the following, we use $\tilde{X}$ to represent perturbations in the synchronous gauge. We set $\tilde{\theta}_{c}$ to zero to fix the synchronous gauge. Then the gauge transformation equations are
$\tilde{\sigma}_{\alpha}$ $=\sigma_{\alpha}$
$\tilde{\theta}_{\alpha}$ $=\theta_{\alpha}-\theta_{c}$
$\tilde{\delta}_{\alpha}+3(1+w_{\alpha})\frac{\cal H}{k^{2}}\tilde{\theta}$ $=\delta_{\alpha}+3(1+w_{\alpha})\frac{\cal H}{k^{2}}\theta$
[…]
$\tilde{h}$ is determined by the condition $\tilde{\theta}_{c}=0$ as
$\tilde{h}^{\prime}$ $=-2\tilde{\delta}_{c}^{\prime}~.$ (B.9)
Here we keep terms up to $\tau^{2}$. Terms that are vanishing up to this order, we keep the the leading non-vanishing term. The AD initial conditions are
$\tilde{h}$ $=$ $-\frac{1}{2}k^{2}\tau^{2}$
$\tilde{\eta}$ $=$ $-1+\frac{5+4R_{\nu}}{12(15+4R_{\nu})}k^{2}\tau^{2}$
$\tilde{\delta}_{\gamma}$ $=$ $\tilde{\delta}_{\nu}=\frac{1}{3}k^{2}\tau^{2}$
$\tilde{\delta}_{c}$ $=$ $\tilde{\delta}_{b}=\frac{1}{4}k^{2}\tau^{2}~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(\textrm{AD})$
$\tilde{\theta}_{b}$ $=$ $\tilde{\theta}_{\gamma}=\frac{1}{36}k^{4}\tau^{3}$~
$\tilde{\theta}_{\nu}$ $=$ $\frac{1}{36}\frac{23+4R_{\nu}}{15+4R_{\nu}}k^{4}\tau^{3}$
$\tilde{\sigma}_{\nu}$ $=$ $\frac{2}{3(15+4R_{\nu})}k^{2}\tau^{2}~.$ (B.10)
The BDI and CDI initial conditions are
$\tilde{h}$ $=$ $\omega_{b}\tau-\frac{3}{8}\omega_{b}\omega\tau^{2}$
$\tilde{\eta}$ $=$ $-\frac{1}{6}\omega_{b}\tau+\frac{1}{16}\omega_{b}\omega\tau^{2}$
$\tilde{\delta}_{\gamma}$ $=$ $\tilde{\delta}_{\nu}=-\frac{2}{3}\omega_{b}\tau+\frac{1}{4}\omega_{b}\omega\tau^{2}$
$\tilde{\delta}_{c}$ $=$ $\tilde{\delta}_{b}-1=-\frac{1}{2}\omega_{b}\tau+\frac{3}{16}\omega_{b}\omega\tau^{2}~~~~~~~~~~~~~~~~~~~~~~(\textrm{BDI})$
$\tilde{\theta}_{b}$ $=$ $\tilde{\theta}_{\gamma}=\tilde{\theta}_{\nu}=-\frac{1}{12}\omega_{b}k^{2}\tau^{2}$
$\tilde{\sigma}_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{b}k^{2}\tau^{3}~,$ (B.11)
$\tilde{h}$ $=$ $\omega_{c}\tau-\frac{3}{8}\omega_{c}\omega\tau^{2}$
$\tilde{\eta}$ $=$ $-\frac{1}{6}\omega_{c}\tau+\frac{1}{16}\omega_{c}\omega\tau^{2}$
$\tilde{\delta}_{\gamma}$ $=$ $\tilde{\delta}_{\nu}=-\frac{2}{3}\omega_{c}\tau+\frac{1}{4}\omega_{c}\omega\tau^{2}$
$\tilde{\delta}_{b}$ $=$ $\tilde{\delta}_{c}-1=-\frac{1}{2}\omega_{c}\tau+\frac{3}{16}\omega_{c}\omega\tau^{2}~~~~~~~~~~~~~~~~~~~~~~(\textrm{CDI})$
$\tilde{\theta}_{b}$ $=$ $\tilde{\theta}_{\gamma}=\tilde{\theta}_{\nu}=-\frac{1}{12}\omega_{c}k^{2}\tau^{2}$
$\tilde{\sigma}_{\nu}$ $=$ $\frac{1}{6(15+2R_{\nu})}\omega_{c}k^{2}\tau^{3}~,$ (B.12)
The NDI is given by
$\tilde{h}$ $=$ $-\frac{3R_{\nu}}{4}\omega\tau+\frac{9R_{\nu}}{32}\omega^{2}\tau^{2}$
$\tilde{\eta}$ $=$ $\frac{R_{\nu}}{8}\omega\tau-\left(\frac{R_{\nu}(1-R_{\nu})}{6(15+4R_{\nu})}k^{2}-\frac{3R_{\nu}}{64}\omega^{2}\right)\tau^{2}$
$\tilde{\delta}_{\gamma}$ $=$ $\frac{4}{3}\tilde{\delta}_{b}=-R_{\nu}+\frac{1}{2}R_{\nu}\omega\tau+\frac{1}{48}R_{\nu}(8k^{2}-9\omega^{2})\tau^{2}$
$\tilde{\delta}_{c}$ $=$ $-\frac{3}{4}R_{\nu}+\frac{3}{8}R_{\nu}\omega\tau-\frac{9}{64}R_{\nu}\omega^{2}\tau^{2}$
$\tilde{\delta}_{\nu}$ $=$ $1-R_{\nu}+\frac{1}{2}R_{\nu}\omega\tau-\frac{1}{48}R_{\nu}(8(1-R_{\nu})k^{2}+9R_{\nu}\omega^{2})\tau^{2}~~~~~~~~~~~~~~~~~~(\textrm{NDI})$
$\tilde{\theta}_{b}$ $=$ $\tilde{\theta}_{\gamma}=-\frac{1}{4}R_{\nu}k^{2}\tau+\frac{R_{\nu}(1+3R_{b}-R_{\nu})}{16(1-R_{\nu})}\omega k^{2}\tau^{2}$
$\tilde{\theta}_{\nu}$ $=$ $\frac{1}{4}(1-R_{\nu})k^{2}\tau+\frac{1}{16}R_{\nu}\omega k^{2}\tau^{2}$~
$\tilde{\sigma}_{\nu}$ $=$ $\frac{(1-R_{\nu})}{2(15+4R_{\nu})}k^{2}\tau^{2}~.$ (B.13)
The DRDI is given by
$\tilde{h}$ $=$ $-\frac{3R_{\rm dr}}{4}\omega\tau+\frac{9R_{\rm dr}}{32}\omega^{2}\tau^{2}$
$\tilde{\eta}$ $=$ $\frac{R_{\rm dr}}{8}\omega\tau-\left(\frac{R_{\rm dr}(1-R_{\rm dr}-R_{\nu})}{6(15+4R_{\rm dr}+4R_{\nu})}k^{2}-\frac{3R_{\rm dr}}{64}\omega^{2}\right)\tau^{2}$
$\tilde{\delta}_{\gamma}$ $=$ $\tilde{\delta}_{\nu}=\frac{4}{3}\tilde{\delta}_{b}=-R_{\rm dr}+\frac{1}{2}R_{\rm dr}\omega\tau+\frac{1}{48}R_{\rm dr}(8k^{2}-9\omega^{2})\tau^{2}$
$\tilde{\delta}_{c}$ $=$ $-\frac{3}{4}R_{\rm dr}+\frac{3}{8}R_{\rm dr}\omega\tau-\frac{9}{64}R_{\rm dr}\omega^{2}\tau^{2}$
$\tilde{\delta}_{\rm dr}$ $=$ $1-R_{\rm dr}+\frac{1}{2}R_{\rm dr}\omega\tau-\frac{1}{48}R_{\rm dr}(8(1-R_{\rm dr})k^{2}+9R_{\rm dr}\omega^{2})\tau^{2}~~~~~~~~~~~~~~~~~~(\textrm{DRDI})$
$\tilde{\theta}_{b}$ $=$ $\tilde{\theta}_{\gamma}=\tilde{\theta}_{\nu}=-\frac{1}{4}R_{\rm dr}k^{2}\tau+\frac{R_{\rm dr}(1+3R_{b}-R_{\rm dr}-R_{\nu})}{16(1-R_{\rm dr}-R_{\nu})}\omega k^{2}\tau^{2}$
$\tilde{\theta}_{\rm dr}$ $=$ $\frac{1}{4}(1-R_{\rm dr})k^{2}\tau+\frac{1}{16}R_{\rm dr}\omega k^{2}\tau^{2}$
$\tilde{\sigma}_{\nu}$ $=$ $-\frac{19R_{\rm dr}}{30(15+4R_{\rm dr}+4R_{\nu})}k^{2}\tau^{2}~,~\tilde{\sigma}_{\rm dr}=\frac{15-15R_{\rm dr}+4R_{\nu}}{30(15+4R_{\rm dr}+4R_{\nu})}k^{2}\tau^{2}~.$ (B.14)